OpenStudy (anonymous):

Roll a standard die (number cube ) once. Find each probability, Find P (not 1,2,3,4,)

OpenStudy (anonymous):

Is the not 1-4 meaning you only nee the P of 5 and 6?

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

right 1,2,3,4, cant be counted!

OpenStudy (anonymous):

P(5 or 6) = 2/6 = 1/3

OpenStudy (anonymous):

so would it be 5 or 6?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

probability is 1/3

OpenStudy (anonymous):

P= 1/3